# Continued radio observations of the persistent radio source associated with FRB20190520B provides insights into its origin

### v1

@@ 3.2 Scintillation as a source of PRS variability
Zhang et al. (2023) analysed temporal variability above 1 GHz and found no significant variation in most bands. A marginal variability was seen at 3 GHz. We extend the same variability analysis, adding our 1.3 GHz observations to the 1.5 GHz observations reported by Zhang et al. (2023). One of the causes of variability of a compact radio source is interstellar scintillation. Scintillation arises due to inhomogeneities in the ionized ISM and can cause apparent flux variations in compact radio sources. We follow the formalism introduced in Narayan (1992) and further explored in Walker (1998) for our analysis here. The scattering strength parameter $\xi$ describes the scintillation due to a screen at a distance $D$ from the observer. $\xi=1$ corresponds to the critical value at which phase changes are substantial, across the characteristic first Fresnel zone $\theta_{F}=\sqrt{c/2\pi\nu D}$, where $\nu$ is the frequency of observation. Now, $\xi$ is expressed as $(\nu_{0}/\nu)^{17/10}$, where $\nu_{0}$ is the transition frequency between the weak scattering ($\xi\ll 1$ and $\nu_{0}<\nu$), and strong scattering ($\xi\gg 1$ and $\nu_{0}>\nu$) regimes. This transitional frequency has been estimated to be $\nu_{0}=12.53$ GHz using the galactic free electron density model in pyNE2001 (Cordes & Lazio, 2002). Given $\nu_{0}=12.53$ GHz, we are in the strong scattering regime as $\nu_{0}>\nu$ (as $\nu=1.5\,GHz$). There are two possible scintillation scenarios: diffractive scintillation and refractive scintillation. Diffractive scintillation is relevant at characteristic frequency bandwidth, of $\xi^{-2}\nu=\nu(\nu_{0}/\nu)^{17/5}$ which is $\sim 1$ MHz for our observations. Our observations span a larger bandwidth, washing out diffractive scintillation. Hence, refractive scintillation is the relevant mechanism here. For, refractive scintillation, the expected modulation index is
@@
Assuming a scattering disk at a distance of $D=1$ kpc from us, the size of the Fresnel zone at $\nu=1.5$ GHz is $\theta_{F}=\sqrt{c/2\pi\nu D}=6.74$. Therefore, the angular size of the screen is
$\theta_{r}=\theta_{F}\xi=276.6\,\rm{mas}$ (5)
@@
$t_{r}=2\left(\frac{\nu_{0}}{\nu}\right)^{11/5}=10.2\,\rm{days}$ (7)
@@
We obtain the observed modulation index $\bm{m_{\rm{obs}}=0.18}$. The observed modulation of the PRS flux is less than that expected from this region of the sky, i.e. $\bm{m_{\rm{obs}}<m_{\rm{exp}}}$. This observation can be interpreted in the following ways:
[…]
  • All the variability observed is due to scintillation, but the pyNE2001 model estimate of the transition frequency is not accurate. If this is the case, we can estimate the transition frequency required to produce the observed modulation. Table 3 lists the expected transition frequency, $\nu_{0}$ which can produce the observed modulation index at $\nu=1.5\,$GHz.
@@
For completeness, we also repeat this analysis after removing the expected decay from the pulsar wind nebula model (Section 3.1). The inferred size and timescale constraints remain broadly consistent in both cases.
Figure 5: Constraint on the radius of the PRS source using the projected size of the screen $R_{r}$. The figure is adapted from Bhandari et al. (2023), after correcting for the radius by dividing all limits by 2.
Table 3: Scintillation analysis and predictions at 1.5 GHz
@@
$\theta_{F}$ ($\mu$as) $R_{r}$ (pc) $R_{s}$ (pc) $\theta_{F}$ ($\mu$as) $R_{r}$ (pc) $R_{s}$ (pc) $\theta_{F}$ ($\mu$as) $R_{r}$ (pc) $R_{s}$ (pc)
Data as it is 10.20 0.29 0.18 21.31 5.17 7.78 6.74 1.63 2.46 2.13 0.52 0.78 15.37 29.12
Removed best fit decay 0.16 7.36 2.33 0.74 14.54 37.59
@@ 3.3 Hypernebula Model
3.4 AGN or IMBH Origin for the Persistent Radio Source
An alternative scenario is that the persistent radio source (PRS) associated with FRB20190520B is physically unrelated to the FRB engine, and instead arises from accretion onto an massive black hole as discussed by Anna-Thomas et al. (2023). In this picture, the PRS represents compact synchrotron emission from a radio-loud, low-Eddington active galactic nucleus (AGN), while the FRB originates from a separate compact object—such as a magnetar—embedded within the same nuclear environment.
Recent work by Dong et al. (2024) has systematically compared the radio spectra of persistent sources in nearby dwarf galaxies—including known low-mass AGNs—with those of FRB-associated PRSs, including FRB20190520B and FRB 20121102A. They find that several AGNs in dwarf hosts exhibit flat or slightly inverted radio spectra and luminosities consistent with both FRB PRSs, suggesting that an AGN origin cannot be ruled out based on spectral properties alone.
[…]
A similar interpretation has been proposed for the PRS associated with FRB 20121102A. Bhardwaj et al. (2025a) argued that its long-term flux stability, compact size ($<0.35$ pc), and placement on the radio-loud fundamental plane of black hole activity can be explained by a low-Eddington AGN powered by an IMBH. Given the morphological and spectral similarities between the two sources, a comparable AGN origin for the PRS of FRB20190520B remains observationally viable. However, as with FRB 20121102A, this model requires that the FRB bursts themselves arise from a distinct, co-located source within the broader accretion environment.
@@ 4 Conclusion
In this paper, we present broadband radio observations (0.3–-10 GHz) of the persistent radio source (PRS) associated with FRB20190520B. We analyze the data under different proposed models and comment on the physics responsible for the observed emission. The findings are summarized below.
@@
  • The first low frequency ($<1\,$GHz) observations of this PRS hint at a spectral break between $\approx 0.7$ GHz and $\approx 1.3$ GHz, which can be attributed to synchrotron self-absorption. It is not possible to estimate the self-absorption frequency using only the upper limits from the low-frequency observations presented here.
  • We fit the decaying light curves at all frequencies using the decay equation described in Equation 3, derived from the magnetar wind nebula model. This resulted in an estimate of the powerlaw index for energy injection $\alpha=1.75^{+0.09}_{-0.08}$ and the age of the magnetar is $t_{\rm{age}}=53^{+17}_{-10}$. These values are in agreement with constraints from previous observations of the PRS (Bhandari et al., 2023; Zhao & Wang, 2021).
  • We also comment on the contribution of scintillation to the variability of the PRS at 1.5 GHz. Through this analysis, a conservative lower limit on the size of the source $>0.52$ pc (for a screen at 10 kpc) is derived. This limit is consistent with constraints from VLBI observations and low-frequency non-detection.
  • Finally, we discuss the observations under the purview of the best-fit hypernebula model for the PRS associated with FRB20190520B presented in (Sridhar & Metzger, 2022). This best-fit model predicts a rise in flux with time at 1.5 GHz, which does not agree with the observations presented here.
@@
We thank the staff of the GMRT that made these observations possible. GMRT is run by the National Centre for Radio Astrophysics of the Tata Institute of Fundamental Research. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. M.B is a McWilliams fellow and an International Astronomical Union Gruber fellow. M.B. also receives support from the McWilliams seed grant.

### v2

@@ 3.3 Scintillation as a source of PRS variability
2023ApJ...959...89Z analysed temporal variability above 1 GHz and found no significant variation in most bands. A marginal variability was seen at 3 GHz. We extend the same variability analysis, adding our 1.3 GHz observations to the 1.5 GHz observations reported by 2023ApJ...959...89Z. One of the causes of variability of a compact radio source is interstellar scintillation. Scintillation arises due to inhomogeneities in the ionized ISM and can cause apparent flux variations in compact radio sources. We follow the formalism introduced in 1992RSPTA.341..151N and further explored in 1998MNRAS.294..307W for our analysis here. The scattering strength parameter $\xi$ describes the scintillation due to a screen at a distance $D$ from the observer. $\xi=1$ corresponds to the critical value at which phase changes are substantial, across the characteristic first Fresnel zone $\theta_{F}=\sqrt{c/2\pi\nu D}$, where $\nu$ is the frequency of observation. Now, $\xi$ is expressed as $(\nu_{0}/\nu)^{17/10}$, where $\nu_{0}$ is the transition frequency between the weak scattering ($\xi\ll 1$ and $\nu_{0}<\nu$), and strong scattering ($\xi\gg 1$ and $\nu_{0}>\nu$) regimes. This transitional frequency has been estimated to be $\nu_{0}=12.53$ GHz using the galactic free electron density model in pyNE2001 (2002astro.ph..7156C). Given $\nu_{0}=12.53$ GHz, we are in the strong scattering regime as $\nu_{0}>\nu$ (as $\nu=1.5$ GHz). There are two possible scintillation scenarios: diffractive scintillation and refractive scintillation. Diffractive scintillation is relevant at characteristic frequency bandwidth, of $\xi^{-2}\nu=\nu(\nu_{0}/\nu)^{17/5}$ which is $\sim 1$ MHz for our observations. Our observations span a larger bandwidth, washing out diffractive scintillation. Hence, refractive scintillation is the relevant mechanism here. For, refractive scintillation, the expected modulation index is
@@
Assuming a scattering disk at a distance of $D=1$ kpc from us, the size of the Fresnel zone at $\nu=1.4$ GHz is $\theta_{F}=\sqrt{c/2\pi\nu D}=6.78\,\mu\rm{as}$. Therefore, the angular size of the screen is
$\theta_{r}=\theta_{F}\xi=284.1\,\mu\rm{as}$ (5)
@@
$t_{r}=2\left(\frac{\nu_{0}}{\nu}\right)^{11/5}=10.47\,\rm{days}$ (7)
@@
We obtain the observed modulation index $\bm{m_{\rm{obs}}=0.23}$. The observed modulation of the PRS flux is less than that expected from this region of the sky, i.e. $\bm{m_{\rm{obs}}<m_{\rm{exp}}}$. This observation can be interpreted in the following ways:
  • All the variability observed is due to scintillation, and the pyNE2001 estimate of the transition frequency is accurate. However, the observed flux modulation is lesser than the expected modulation. This change in modulation can be explained if the size of the source $\theta_{s}$ is slightly larger than the size of the screen $\theta_{r}$, reducing the modulation index by a factor of $(\theta_{r}/\theta_{s})^{7/6}$, and increasing refractive timescale by a factor of $\theta_{s}/\theta_{r}$ (see 1998MNRAS.294..307W, for details). Table 3 summarizes the possible constraints on the size of the source ($R_{s}=\theta_{s}d_{\rm{lum}}/2$), and the modified timescale of the scintillation, for screens at a distance of 0.1 kpc, 1 kpc and 10 kpc. The projected screen radius ($R_{r}$) values listed in Table 3 can be quoted as a conservative lower limit of the size of the source. These are shown in Figure 4 for different distances of the screen from the observer in the $t_{\rm{age}}$$R_{n}$ phase space plot, adapted from 2023ApJ...958L..19B.
  • All the variability observed is due to scintillation, but the pyNE2001 model estimate of the transition frequency is not accurate. If this is the case, we can estimate the transition frequency required to produce the observed modulation. Table 3 lists the expected transition frequency, $\nu_{0}$ which can produce the observed modulation index at $\nu=1.4\,$GHz.
@@
For completeness, we also repeat this analysis after removing the expected decay from the magnetar wind nebula model (Section 3.2). The inferred size and timescale constraints remain broadly consistent in both cases.
Figure 4: Constraint on the radius of the PRS source using the projected size of the screen $R_{r}$. The figure is adapted from 2023ApJ...958L..19B, after correcting for the radius by dividing all limits by 2.
Table 3: Scintillation analysis and predictions at $\approx$1.4 GHz (the observation frequency used is a mean of all the detections from Table 1 i.e. $\nu\approx 1.4$ GHz)
@@
[…]
Data as it is 10.47 0.29 0.23 21.44 5.30 6.48 6.78 1.68 2.05 2.14 0.53 0.65 12.80 18.93
Removed best fit decay 0.22 6.80 2.15 0.68 13.43 20.92
@@ 3.4 Hypernebula Model
Another model proposed to explain the PRSs associated with FRBs is the hypernebula model (2022ApJ...937....5S). In this model, mass transfer from a companion star onto a massive black hole via Roche lobe overflow drives powerful disk winds. These winds inflate a bubble of relativistic electrons that emit synchrotron radiation. Two key timescales characterize the model: the timescale over which the wind expands freely into the CSM, $t_{\rm{free}}$; and $t_{\rm{active}}\sim 10$$10^{6}$ years, the total duration of the accretion phase (2022ApJ...937....5S). The relatively modest evolution in burst DM over time (2022Natur.606..873N; 2023Sci...380..599A) suggests that the system is currently at $t<t_{\rm{free}}$. However, 2021ApJ...923L..17Z show a decreasing trend in the burst DM. (2022ApJ...937....5S) applied the hypernebula model specifically to FRB20190520B, using a set of best-fit model parameters (e.g., age, mass accretion rate, ambient density) chosen to match the source’s known properties at the time. For those specific parameters, their model predicts a gradual increase in radio flux of the PRS at GHz frequencies over decadal timescales—this is explicitly stated in Section 4.2 of their paper and shown in their model curves (see Figures 5 and 6 in 2022ApJ...937....5S). Whereas, we see a slow decay in the flux over the span of $\sim 4$ years. This is in contrast to our observations, indicating that the best-fit hypernebula model using data from the first 2 years since detection is inconsistent with the recent observed behavior of the PRS associated with FRB20190520B.
3.5 AGN or IMBH Origin for the Persistent Radio Source
An alternative scenario is that the persistent radio source (PRS) associated with FRB20190520B is physically unrelated to the FRB engine, and instead arises from accretion onto a massive black hole as discussed by 2023Sci...380..599A. In this picture, the PRS represents compact synchrotron emission from a radio-loud, low-Eddington active galactic nucleus (AGN), while the FRB originates from a separate compact object—such as a magnetar—embedded within the same nuclear environment.
Recent work by 2024ApJ...973..133D has systematically compared the radio spectra of persistent sources in nearby dwarf galaxies—including known low-mass AGNs—with those of FRB-associated PRSs, including FRB20190520B and FRB 20121102A. They find that several AGNs in dwarf hosts exhibit flat or slightly inverted radio spectra and luminosities consistent with both FRB PRSs, suggesting that an AGN origin cannot be ruled out based on spectral properties alone.
@@ 3.5 AGN or IMBH Origin for the Persistent Radio Source
A similar interpretation has been proposed for the PRS associated with FRB 20121102A. 2025PASP..137h4202B argued that its long-term flux stability, compact size (<0.35<0.35 pc), and placement on the radio-loud fundamental plane of black hole activity can be explained by a low-Eddington AGN powered by an IMBH. Given the morphological and spectral similarities between the two sources, a comparable AGN origin for the PRS of FRB20190520B remains observationally viable. However, as with FRB 20121102A, this model requires that the FRB bursts themselves arise from a distinct, co-located source within the broader accretion environment.
@@ 4 Conclusion
In this paper, we present broadband radio observations (0.30.3–-$10$ GHz) of the persistent radio source (PRS) associated with FRB20190520B. We analyze the data under different proposed models and comment on the physics responsible for the observed emission. The findings are summarized below.
@@
  • The first low frequency (<1<1\,GHz) observations of this PRS hint at a spectral break between $\approx 0.7$ GHz and $\approx 1.3$ GHz, which can be attributed to synchrotron self-absorption. It is not possible to estimate the self-absorption frequency using only the upper limits from the low-frequency observations presented here.
  • We fit the decaying light curves at all frequencies using the decay equation described in Equation 3, derived from the magnetar wind nebula model. This resulted in an estimate of the powerlaw index for energy injection $\alpha=1.77^{+0.07}_{-0.08}$ and the age of the magnetar is $t_{\rm{age}}$, is $52^{+16}_{-10}$ years. These values are in agreement with constraints from previous observations of the PRS (2023ApJ...958L..19B; 2021ApJ...923L..17Z; 2024arXiv241219358B).
  • We also comment on the contribution of scintillation to the variability of the PRS at 1.5 GHz. Through this analysis, a conservative lower limit on the size of the source >0.53>0.53 pc (for a screen at 10 kpc) is derived (assuming that NE2001 accurately models the transition frequency and that the observed variability is due to scintillation). This limit is consistent with constraints from VLBI observations and low-frequency non-detection. We also note that the variability could be a combination of scintillation and intrinsic variability, and it is not trivial to distinguish between the two.
  • Finally, we discuss the observations under the purview of the best-fit hypernebula model for the PRS associated with FRB20190520B presented in (2022ApJ...937....5S). This best-fit model predicts a rise in flux with time at 1.5 GHz, which does not agree with the observations presented here.
@@
[…]
Appendix A Comparison of calibration of images
This section contains a comparison of the image of the target PRS along with the calibrator, 3C286 to demonstrate the calibrations errors present in the 2022 November 11 data. This is shown in Figure 5.
Figure 5: Comparison of the image of the target PRS (top row) and the calibrator 3C286 (bottom row) on 11 Nov 2022 (left column) and 29 Nov 2022 (right column). This shows the calibration errors in the 11 Nov 2022 data.